Solve: xy=1 and 9y=x-7
Must I use the quadratic equation? I can't seem to come up with the right answer even when I use it.

Answer by mananth(16946)

xy=1 and 9y=x-7
xy=1
so y= 1/x
plug the value y in the equation 9y=x-7
9*1/x = x-7
9= x(x-7)
9=x^2-7x
x^2-7x-9=0
x1= 7+sqrt(49+36)/ 2
x1= 8.1
x2=7-sqrt(49+36)/ 2
x2=-1.12
..
Plug the value of x1 in xy=1
y1=1/8.1
y1=0.12
..
y2 = 1/-1.1= -0.90
y2= 0.1234
..
xy has two values
(-1.12, -0.9) & (8.1 , 0.1234)